Economic indicators before taxes

Net Present value (NPV) ($000) $664,451

IRR % 21.2%

payback (years) 3.24

Discount rate 5%

Economic indicators after taxes

NPV ($000) $ 396,887

IRR % 15.8%

Payback(years) 4.10

discount rate 5%

what happens to the NPV if the discount rate is increased to 10,15 and 30% respectively?
Answer #1

Let x be the cash inflows flow, y be the cash outflow per year

Let number of years be 5 years

By data ,Net present value = 664451

Present value of net cash inflows- Present value of cash outflows = 664451

4.3295x -y = 664451 .......... Equation-1

By data, IRR =21.2%

2.9133x = y......................Equation-2

Solving equation-`1 and 2 we get'

X = 469,199 , y=13,66,933

NPV WHEN DISCOUNT RATE IS 10%

NPV = (3.7908*469199) - 1366933 = 411700

NPV when discount rate is 15 %

NPV = (3.3521*469199) - 1366933 = 205895

NPV when discount is 30%

NPV = (2.4356*469199) - 1366933 = -224166

Note - When the assumptions to the question the answer changes

- Annuity values were considered for computation purposes